Flight Café prepares in-flight meals for airlines in its kitchen located next to a local airport. The companys planning budgIn July, 27,000 meals were actually served. The companys flexible budget for this level of activity appears below: Flight CaRequired: 1. Calculate the companys activity variances for July. (Indicate the effect of each variance by selecting F for
Answer #1

Answer

[1]

Flight Café
Activity Variance = Difference between Flexible budget and Planning Budget
Revenue $4,400 Favourable
Expenses:
Raw materials $2,100 Unfavourable
Wages & Salaries $200 Unfavourable
Utilities $50 Unfavourable
Facility Rent $0 None
Insurance $0 None
Miscellaneous $100 Unfavourable
Total Expenses $2,450 Unfavourable
Net Operating Income $1,950 Favourable

--Working

Flight Café
Activity Variance = Difference between Flexible budget and Planning Budget
Revenue =118800-114400 Favourable
Expenses:
Raw materials =56700-54600 Unfavourable
Wages & Salaries =11400-11200 Unfavourable
Utilities =3150-3100 Unfavourable
Facility Rent =3000-3000 None
Insurance =2900-2900 None
Miscellaneous =3200-3100 Unfavourable